If a,b,c......in A.P so proved that (b+c),(c+a),(a+b)......in A.P Plz give me solution

Grade:12th pass

2 Answers

prateek
7 years ago
FOR A.P. a+c=2b
and 2(a+c)=(b+c) + (a+b)
      2(a+c)=b+c+a+b
      2(a+c)=a+c+2b
      2(a+c)=a+c + a+c     {2b=a+c}
      2(a+c)=2(a+c)
I HOPE YOU WILL FIND MY ANSWER HELPFUL THANKS
PLEASE CHECKOUT THE COURSE THAT WE PROVIDE.
PLZ POST ANY QUERIES THAT U HAVE
      
Ankur Saha
7 years ago
as a,b,c are in A.P. so 2b=a+c..................(i)
If (b+c) , (a+c) , (a+b) will be in A.P. ,,then (b+c) + (a+b)=2(a+c).
Now (b+c) + (a+b)= b+c+a+b =2b+a+c =2b+2b    {from(i)}     
                                                           =4b =2.2b =2(a+c)   {from (i)}
SO,,,,,,(b+c) , (a+c) , (a+b) are in A.P.                 [PROVED]